OpenStudy (bobobox):

I need some help I need to get atleast over 70% on this assignment or else I can't take my final exams. It's five questions about the division of monomials.

OpenStudy (bobobox):

@TheSmartOne

OpenStudy (bobobox):

First question.

OpenStudy (anonymous):

square the top what do you get?

OpenStudy (bobobox):

16x

OpenStudy (anonymous):

hmm no \[(4x)^2=4x\times 4x\] which is not \(16x\)

OpenStudy (bobobox):

this is hard Can you explain to me why? @TheSmartOne

OpenStudy (anonymous):

4x times 4x = 4x^2 you forgot to multiply the variables

OpenStudy (anonymous):

I MEAN 16x^2 sorry

OpenStudy (bobobox):

okay, so would it be 

OpenStudy (bobobox):

then which answer would it be? I think its d

OpenStudy (anonymous):

no

OpenStudy (bobobox):

c?

OpenStudy (anonymous):

when multiplying fractions you multiply the denominators 4x/y * 4x/y = 16x^2/y^2

OpenStudy (bobobox):

oh yes thats right you cant just multiply the numerator @090909090909

OpenStudy (bobobox):

So c right? could you possibly help me with the next?

OpenStudy (anonymous):

yah. try the next ones first though and I'll check them

OpenStudy (bobobox):

okay! one sec

OpenStudy (bobobox):